OpenStudy (anonymous):

find the slope- intercept equation for the line with the indicated slope and y-intercept SLOPE\[\frac{ 4 }{ 7 }\]; y intercept (0,1)

OpenStudy (anonymous):

y= 4/7x +1 y=mx + b m= slope which is 4/7, (0,1) if you have a 0 where the x value is the y intercept would be 1 since b represents the horizontal slope I think... lol I cant explain it right but i can guarantee thats the answer.